### v3

@@ 4.3 SC effectively identifies the underlying LDSs
In this subsection, we compare the prediction accuracy of the identified LDSs by each approach. We first present the average reconstruction error across all states at each time step for the UCSD dataset in Fig. 3. For illustrative purposes, we select $r=3,20$ and $30$ as examples. It is clearly shown that as the input dimension increases (e.g., $r=30$), LS fails to reconstruct the input states. However, SC consistently achieves results comparable to CG, WLS, and SOC, while being orders of magnitude faster for computation and highly efficient in memory usage (see Fig. 1 (a) and Fig. 1 (c)).
We then show the average reconstruction error for the DTDB dataset in Fig. 3. Given the varying time steps of the input sequences, the reconstruction error is averaged over the time percentiles. The results indicate that SC outperforms SOC since it achieves a lower reconstruction error. Note that though the LS approach achieves the lowest error within the training data horizon, it does not guarantee stability, which is crucial for long-term prediction (see Section. 4.4).
@@
Figure 3: Average reconstruction error evaluated at each time step percentile of the DTDB dataset.
4.4 SC enables stable long-term predictions
@@
We first select two example sequences (i.e., flushing toilet and running stream) and show the qualitative results in Fig. 6. These results indicate that LS approach mostly learns the LDSs that blow up during the extended horizon, as it does not provide with stability guarantee. Conversely, the SOC approach tends to learn LDSs that converge to zeros, indicating a lack of dynamic behavior.
@@
Magnitude: i) Turning black: if the pixels of the final frame are all zero, ii) Blowing up: if the pixels of the final frame blow up, and iii) Normal: neither Turning black nor Blowing up. These characterizations aim to determine if the system evolution of the LDSs is still meaningful.
Dynamics: i) Moving: if the last few frames still exhibit the dynamical pattern. This characterization aims to determine if the system evolution of the LDSs still reflects the dynamic pattern.
Turning black ($\downarrow$) Blowing up ($\downarrow$) Normal ($\uparrow$) Moving ($\uparrow$)
LS 6.67% 45.96% 47.37% 62.11%
SOC 23.16% 0% 76.84% 22.11%
SC 14.39% 0% 85.61% 54.39%
Table 1: Quantitative metrics (ratios) characterizing predictions from each method.
In Table. 1, we report the quantitative ratios in terms of the defined metrics over the 285 sequences in the DTDB dataset. Note that the values for LS method are marked in gray as this method does not guarantee stability. The results indicate that both SOC and SC learn stable LDSs without blow-ups thanks to the stability guarantee, but SOC is less effective at modeling long-term dynamic patterns, as the learned LDSs often converge to equilibrium states (i.e., not moving) or even complete zeros (i.e., turning black). On the other hand, although the LDSs learned by the LS approach exhibit the most dynamic patterns, they tend to be unstable, leading to blow-ups.
@@ 4.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
In this subsection, we demonstrate that when combined with the Koopman Operator theory (Koopman, 1931), SC can also be effectively and efficiently applied to highly nonlinear robot manipulation tasks. For this, we introduce a highly nonlinear robot manipulation dataset.
DexManip (Han et al., 2023; Rajeswaran et al., 2018b): A dataset of Adroit Hand performing two dexterous manipulation tasks in MUJOCO simulation environment (Todorov et al., 2012). Though the original state evolution is nonlinear, but as shown in Han et al. (2023), via lifted to a higher-dimension Hilbert space (e.g., $r\approx 750$), the system can be considered linear in the lifted state space. Two tasks are briefly described as follows: i) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and ii) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. A for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
@@
We first report the computation time and memory usage in Table. 2. As seen, SC also alleviates the computational and memory burden compared to the SOC method. Note that the values for LS method are marked in gray as this method does not guarantee stability. We then show the average reconstruction error on the DexManip datasets in Fig. 7 (a) & (b), where SC demonstrates the best performance, while LS results in significantly larger errors on the Tool dataset due to its lack of stability.
@@
To further underscore the importance of ensuring stability when learning LDSs for robotic dexterous manipulation tasks in the DexManip datasets, we also compute the task success rates and the safety rates (both reported in Table. 2) for each task when tracking the generated robot hand trajectories. The definition of these two metrics are included in Appendix. A.
For both tasks, SC learns stable LDSs that achieve the highest task success rates and complete the tasks safely. In contrast, LS fails to complete tasks safely, while SOC underperforms due to the fact it learns overly conservative LDSs that lack the ability to generate diverse trajectories.
@@ 4.6 SC can learn stable LDSs from shorter and corrupted datasets
In real-world settings, the collected demonstrations may be imperfect, unlike those in the original DexManip datasets. Hence, we introduce two additional datasets: (i) Tool_short, where each demonstration is shorter, posing challenges in learning stable LDSs with a limited horizon, and (ii) Reloc_with_failures, which includes a few failed demonstrations (see Appendix. A for more details).
@@
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in Rowley et al. (2009), we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset.
Specifically, for each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of system matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following (Rowley et al., 2009), we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from $Reloc\_with\_failures$ dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 8, respectively. It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 3). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 8. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 3).
Figure 8: Rollout with two subsets of the Koopman modes. The green box represents rollouts using Unstable Set (norm of eigenvalue larger than 1). The blue box represents rollout with the rest Stable Set. In the green box, the upper line shows the rollout before clipping and the lower line shows it after clipping. It is obvious that after clipping, the rollout becomes stable, thereby allowing the composite rollout of both sets to safely complete the task (as shown in Table 3).
@@ 4.7 SC can identify systems with control inputs
In this subsection, we demonstrate the comparable capability of SC in identifying stable LDSs with control inputs (see Appendix. B for details). To this end, we first introduce another dataset that includes the control commands.
Franka Panda (Gaz et al., 2019): A dataset of Franka Panda Robot Arm’s movements. The robot state space has 17 dimensions ($r=17$), including the coordinates of end effector ($\mathbb{R}^{3}$), joint angle values ($\mathbb{R}^{7}$), and joint angular velocities ($\mathbb{R}^{7}$), and the control input has $\mathbb{R}^{7}$ joint torques. This dataset consists of 3100 time steps of control inputs (randomly generated) and the corresponding robot states.
@@
We first evaluate the long-term prediction accuracy of each approach. For this, we first command the random control inputs in simulation with ground-truth system dynamics to generate test trajectories. We then rollout the trained $A$ and $B$ matrices with the same control inputs to compute the Reconstruction error for each approach. From the results shown in Fig. 9, we can see that SC stills perform comparably to all other approaches.
Additionally, we use the trained $A$ and $B$ matrices to construct an LQR controller, and use such controller to track a figure-8 pattern in simulation (similar in Mamakoukas et al. (2020)). How accurately the figure is tracked indicates how well the learned LDS aligns with the actual dynamical system with control inputs. For each approach, we repeat this process five times with different initial positions of the end effector. We show the qualitative tracking performances in Appendix. C.
Figure 9: Distance to correct end effector position calculated at each time step. Solid lines represent median values, and shaded areas indicate quartile values over five policy rollouts.
@@ 5 Limitations and Future Work
Though spectrum clipping demonstrates its efficacy on identifying the stable LDSs over various datasets, there are still opportunities for future study. First, the performance on the dataset with control inputs is not yet fully comparable to other state-of-the-art approaches, indicating potential for algorithmic improvements. Second, while we demonstrate the effectiveness of SC experimentally, a deeper analysis of the spectrum properties before and after clipping could provide valuable theoretical insights. Third, if the system matrix is not diagonalizable, spectrum clipping may not be applicable. For example, if $A_{ls}=[1,1,0;0,1,1;0,1,1]$, the decomposition and recomposition algorithm introduced in Section. 3 can not maintain the original matrix. However, our experiments showed that a tiny Gaussian noise $N(0,10^{-24})$ added to each element of this matrix helped maintaining its structure after de- and re-composition. Therefore, this may not be a concern for the applications considered in this work, as learned matrices typically do not have a noise-free structure.
@@ 6 Conclusion
Appendix A Details of DexManip tasks
DexManip is a dataset collected when using Adroit Hand to conduct two different dexterous manipulation tasks in MUJOCO simulation environment. In this dataset, the states of hand and states of object is captured, concatenated and lifted to a higher-dimension Hilbert space at each time step (see Han et al. (2023)), so that the hand-object system becomes an LDS in the Hilbert space. Thereafter, we use a simple PD controller to track the generated hand trajectories in the simulator. The tasks are detailed as follows.

### v4

@@ 4.2 SC significantly reduces computation & memory burden
LS CG [2] WLS [24] SOC [26, 27] SC
Time complexity -- $O(n^{6})$ $O(n^{6})$ $O(n^{3})$ $O(n^{3})$
Space complexity $O(n^{2})$ $O(n^{4})$ $O(n^{4})$ $O(n^{2})$ $O(n^{2})$
Table 1: Space and Time complexity of each method. $n$ is the system dimension. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. For iterative methods (CG, WLS, and SOC), the shown time complexity corresponds to a single iteration.
In addition to empirical results on the two datasets, we report the time and space complexity of each method in Table. 1, further supporting the observation that SC is as or more efficient in both time and memory than the other methods. For time complexity, since all methods modify the original LS solution to enforce stability, we compare only the cost incurred after solving LS. Note that CG, WLS, and SOC are iterative algorithms, and the table reports the time complexity per iteration. In contrast, SC is non-iterative, and its overall time complexity is shown.
@@ 4.3 SC effectively identifies the underlying LDSs
In this subsection, we compare the prediction accuracy of the identified LDSs by each approach. We first present the average reconstruction error across all states at each time step for the UCSD dataset in Fig. 2 (left). For illustrative purposes, we select $r=3,20$ and $30$ as examples. It is clearly shown that as the input dimension increases (e.g., $r=30$), LS fails to reconstruct the input states, with the error increasing significantly over time. A similar trend can be observed in Fig. 3, where we show the average reconstruction error over all states and time steps of each method at each training dimension. In both cases, SC consistently achieves results comparable to CG, WLS, and SOC, while being orders of magnitude faster for computation and highly efficient in memory usage (see Section. 4.2).
We then show the average reconstruction error for the DTDB dataset in Fig. 2 (right). Note that due to the high dimension of the state space in this dataset, only LS, SOC, and SC are applicable. Both CG and WLS involve solving a Kronecker product, which has a significantly large space complexity (i.e., $O(n^{4})$ shown in Table. 1, $n$ is the size of the state space), leading to memory overflow.
Given the varying time steps of the input sequences, the reconstruction error is averaged over the time percentiles. The results indicate that SC outperforms SOC since it achieves a lower reconstruction error. Note that though the LS approach achieves the lowest error within the training data horizon, it does not guarantee stability, which is crucial for long-term prediction (see Section. 4.4).
Figure 2: Left figure: average reconstruction error evaluated on different subspace dimensions of the UCSD dataset. Shaded areas indicate quartile values over all sequences. Right figure: average reconstruction error evaluated at each time step percentile of the DTDB dataset. Shaded areas indicate quartile values over all sequences.
4.4 SC enables stable long-term predictions
@@
We first select two example sequences (i.e., flushing toilet and running stream) and show the qualitative results in Fig. 6. These results indicate that LS approach mostly learns the LDSs that blow up during the extended horizon, as it does not provide with stability guarantee. Conversely, the SOC approach tends to learn LDSs that converge to zeros, indicating a lack of dynamic behavior. In Table. 2, we report the spectral radius of the learned system matrix for each method. The results support the conclusion that LS leads to blow-ups due to a spectral radius greater than one, while SOC results in vanishing values as its spectral radius remains strictly less than one.
LS SOC [26, 27] SC
DTDB Sequence 185 1.0025 0.9999 1.0000
DTDB Sequence 205 1.0022 0.9999 1.0000
Table 2: Spectrum radii of the system matrices learned with each method.
@@
Quality (NIQE score [44]): A commonly used image quality evaluation metric. This metrics helps determine if the predictions of the LDSs are still meaningful. Note that images with extremely low quality (for example, when most pixel values converge to 0 or blow up) can produce unbounded NIQE scores. We substitute those unbounded scores with $10^{5}$ to provide a bounded statistical summaries.
Dynamics (Moving Ratio  [^1] Moving is defined as $||F_{M}-F_{M-2}||_{1}>T$, where $\{F_{i}\}_{i=1}^{M}$ is a sequence of $M$ image frames and $T=9$.): The ratio of sequences in which the final frames still exhibit the dynamic pattern This characterization aims to determine if the system evolution of the LDSs still reflects the dynamic pattern.
Average NIQE ($F=500$) ($\downarrow$) Average NIQE ($F=3000$) ($\downarrow$) Moving ($\uparrow$)
LS 904.6265 6355.5817 62.11%
SOC [26, 27] 832.8063 4243.4279 22.11%
SC 18.8769 1075.0140 54.39%
[…]
Table 3: Quantitative metrics evaluating the predictions of each method on the DTDB dataset.
In Table. 3, we report the quantitative ratios in terms of the defined metrics over the 285 sequences in the DTDB dataset. Note that the values for LS method are marked in gray as this method does not guarantee stability. The results indicate that SC best preserves long-term dynamic patterns while maintaining high prediction accuracy, as evidenced by its superior image quality—particularly at 500-step predictions, which are nearly comparable to the training set images.
@@ 4.5 SC’s benefits extend to nonlinear systems when combined with the Koopman Operator
In this subsection, we demonstrate that when combined with the Koopman Operator theory [7], SC can also be effectively and efficiently applied to highly nonlinear robot manipulation tasks. For this, we introduce a highly nonlinear robot manipulation dataset.
DexManip [13, 45]: A dataset of Adroit Hand performing two dexterous manipulation tasks in MUJOCO simulation environment [46]. Though the original state evolution is nonlinear, but as shown in [13], via lifted to a higher-dimension Hilbert space (e.g., $r\approx 750$), the system can be considered linear in the lifted state space. Two tasks are briefly described as follows: i) Tool Use: The robot hand is tasked with picking up the hammer and driving the nail into the board placed at a random height, and ii) Object Relocation: The robot hand is tasked with moving a cylinder to a random target location (see Appendix. B for the details of each dataset). Similar to the DTDB dataset, only LS, SOC, and SC are applicable due to the high dimensionality of the lifting space.
@@
We first report the computation time and memory usage in Table. 4. As seen, SC also alleviates the computational and memory burden compared to the SOC method. Note that the values for LS method are marked in gray as this method does not guarantee stability. We then show the average reconstruction error on the DexManip datasets in Fig. 7 (a) & (b), where SC demonstrates the best performance and LS results in significantly larger errors on the Tool dataset due to its lack of stability.
@@
To further underscore the importance of ensuring stability when learning LDSs for robotic dexterous manipulation tasks in the DexManip datasets, we also compute the task success rates and the safety rates (both reported in Table. 4) for each task when tracking the generated robot hand trajectories. The definition of these two metrics are included in Appendix. B.
For both tasks, SC learns stable LDSs that achieve the highest task success rates and complete the tasks safely. In contrast, SOC underperforms due to the fact it learns overly conservative LDSs that lack the ability to generate diverse trajectories. Note that while LS seems to achieve the desired performance on the Reloc task, its effectiveness drops significantly when trained on a corrupted dataset as shown in Section. 4.6.
@@ 4.6 SC can learn stable LDSs from shorter and corrupted datasets
In real-world settings, the collected demonstrations may be imperfect, unlike those in the original DexManip datasets. Hence, we introduce two additional datasets: (i) Tool_short, where each demonstration is shorter, posing challenges in learning stable LDSs with a limited horizon, and (ii) Reloc_with_failures, which includes a few failed demonstrations (see Appendix. B for more details).
@@
We also reveal the connections between SC and the Koopman Operator’s spectral properties. Following the algorithm described in [33], we can rollout the systems with only subsets of the Koopman modes and visualize the hand motions to investigate the effectiveness of SC upon each subset.
Specifically, for each eigenvalue $\lambda_{i}\ (i=1,...,n)$ of Koopman matrix $K$, we can find corresponding eigenvector $\mathbf{v_{i}}$ (i.e., the mode) and eigenvector $\mathbf{w_{i}}$ of adjoint matrix $K^{*}$, such that $K\mathbf{v_{i}}=\lambda_{i}\mathbf{v_{i}}\text{, and }K^{*}\mathbf{w_{i}}=\overline{\lambda_{i}}\mathbf{w_{i}}$. Then following [33], we construct the corresponding eigenfunction $\varphi_{i}:\mathbb{R}^{n}\rightarrow C$, defined as $\varphi_{i}(\mathbf{z})=\langle\mathbf{z},\mathbf{w}_{i}\rangle$. By doing so, the rollout of eigenfunctions can be represented by $\varphi_{i}(K\mathbf{z})=\lambda_{i}\varphi_{i}(\mathbf{z})$, and the states can be reconstructed by $\mathbf{z_{k}}=\sum_{i=1}^{n}\lambda_{i}^{k}\varphi_{i}(\mathbf{z_{0}})\mathbf{v_{i}}$. Based on this formula, specific modes can be selected for rollouts.
In our experiments, we first computed the Koopman matrix $K_{f}$ from $Reloc\_with\_failures$ dataset. We then sorted the eigenvalues by their norms from largest to smallest and divided them into two subsets for rollouts: (i) Unstable Set: the first 52 Koopman modes with unstable eigenvalues, which have norms greater than 1, and (ii) Stable Set: the remaining 707 modes. We then visualized the robot hand motions of each subset via the eigenfunction rollouts in Fig. 8, respectively (see the supplementary video for full policy rollouts). It can be seen that the rollout of the Unstable Set before clipping clearly generate unstable hand motions (see upper line of the green box), which results in the low success rate and safety rate when combined with Stable Set for the final task performance (as shown in Table. 5). Further, we clipped the matrix $K_{f}$ and visualized the rollouts of the clipped Unstable Set in Fig. 8. It is obvious that after clipping, the robot hand motions become stable (see bottom line of the green box), enabling the composite system to safely complete the task (as also shown in Table 5). The full policy rollout is also available in the supplementary video.
Figure 8: Visualizations of rollout with two subsets of the Koopman modes. The green box represents rollouts using Unstable Set (norm of eigenvalue larger than 1). The blue box represents rollout with the rest Stable Set. In the green box, the upper line shows the rollout before clipping and the lower line shows it after clipping. It is obvious that after clipping, the rollout becomes stable, thereby allowing the composite rollout of both sets to safely complete the task (as shown in Table 5. The full policy rollout is available in the supplementary video.
@@ 4.7 SC can identify systems with control inputs
In this subsection, we demonstrate the comparable capability of SC in identifying stable LDSs with control inputs (see Appendix. C for details). To this end, we first introduce another dataset that includes the control commands.
Franka Panda [35]: A dataset of Franka Panda Robot Arm’s movements. The robot state space has 17 dimensions ($r=17$), including the coordinates of end effector ($\mathbb{R}^{3}$), joint angle values ($\mathbb{R}^{7}$), and joint angular velocities ($\mathbb{R}^{7}$), and the control input has $\mathbb{R}^{7}$ joint torques. This dataset consists of 3100 time steps of control inputs (randomly generated) and the corresponding robot states.
@@
We first evaluate the long-term prediction accuracy of each approach. For this, we first command the random control inputs in simulation with ground-truth system dynamics to generate test trajectories. We then rollout the trained $A$ and $B$ matrices with the same control inputs, and then compute the average $L2$ distance between the predicted and ground-truth end-effector positions for each method. From the results shown in Table. 6, We observe that while SC performs worse than CG and SOC when $d=100$, its performance becomes comparable as $d$ increases. Notably, SC achieves this with significantly lower computation time and memory usage.
Additionally, we use the trained $A$ and $B$ matrices to construct an LQR controller, and use such controller to track a figure-8 pattern in simulation (similar in [26]). How accurately the figure is tracked indicates how well the learned LDS aligns with the actual dynamical system with control inputs. For each approach, we repeat this process five times with different initial positions of the end effector. We show the qualitative tracking performances in Appendix. D, where SC demonstrates performance comparable to that of other methods.
LS CG [2] WLS [24] SOC [26, 27] SC
$d=100$ $0.0748\pm 0.1218$ $0.0166\pm 0.0066$ $0.3390\pm 0.2059$ $0.0225\pm 0.0131$ $0.0693\pm 0.1370$
$d=2000$ $0.0207\pm 0.0173$ $0.0154\pm 0.0058$ $0.3051\pm 0.1853$ $0.0154\pm 0.0058$ $0.0183\pm 0.0138$
$d=3000$ $0.0164\pm 0.0083$ $0.0155\pm 0.0058$ $0.2845\pm 0.1501$ $0.0155\pm 0.0058$ $0.0156\pm 0.0057$
Table 6: Average prediction error of the end-effector’s position (unit: $m$).
@@ 5 Limitations and Future Work
Though spectrum clipping (SC) demonstrates its efficacy on identifying the stable LDSs over various datasets, there are still opportunities for future study. First, the performance on the dataset with control inputs is not yet fully comparable to other state-of-the-art approaches, indicating potential for algorithmic improvements. Second, while we empirically demonstrate the effectiveness of SC, a deeper theoretical analysis will likely offer valuable insights.
@@ 6 Conclusion
Appendix A Proof for Arbitrarily Small System Evolution Error
Theorem A.1.
For any $\eta>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, we can always find its diagonalizable counterpart $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$.
Proof.
$\tilde{A}_{LS}$ can be denoted as
$\tilde{A}_{LS}=\hat{A}_{LS}+E$ (3)
where $E\in\mathbb{R}^{n\times n}$ is the residual matrix and, from Lemma. 3.1, $||E||_{1}$ can be arbitrarily small (i.e., $||E||_{1}<\gamma$). Then we have
$\tilde{A}_{LS}^{k}\mathbf{z}=(\hat{A}_{LS}+E)^{k}\mathbf{z}=\hat{A}_{LS}^{k}\mathbf{z}+\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}$ (4)
Therefore,
$\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}=-\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}$ (5)
Using triangle inequality, we have
$||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}=||\sum_{i=1}^{k}{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}||_{1}<=\sum_{i=1}^{k}||{\binom{k}{i}\hat{A}_{LS}^{k-i}E^{i}\mathbf{z}}||_{1}$ (6)
Since $\hat{A}_{LS}$, $k$ and $\mathbf{z}$ are all bounded, there exists a bounded $m>0$, s.t.
$||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<m\gamma$ (7)
Let $\eta=m\gamma$. Note that since Lemma. 3.1 guarantees the existence of a diagonalizable matrix $\tilde{A}_{LS}$ that satisfies an arbitrarily small $\gamma$, it follows that $\eta=m\gamma$ can also be made arbitrarily small. Therefore, for any $\eta>0$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. for any $\mathbf{z}\in\mathbb{R}^{n}$ and any bounded $k\in\mathbb{N}$, $||\hat{A}_{LS}^{k}\mathbf{z}-\tilde{A}_{LS}^{k}\mathbf{z}||_{1}<\eta$. ∎
Appendix B Details of DexManip tasks
DexManip is a dataset collected when using Adroit Hand to conduct two different dexterous manipulation tasks in MUJOCO simulation environment. In this dataset, the states of hand and states of object is captured, concatenated and lifted to a higher-dimension Hilbert space at each time step (see [13]), so that the hand-object system becomes an LDS in the Hilbert space. Thereafter, we use a simple PD controller to track the generated hand trajectories in the simulator. The tasks are detailed as follows.